An automobile accelerates uniformly from rest and reaches a velocity of 22 m/s in 9 s. The tire 

diameter is 58 cm. Determine the number of revolutions the tire makes during this motion 

and the final angular velocity of the tire in revolutions per second


1
Expert's answer
2021-03-16T08:39:48-0400

Given:

"\omega_0=0\: \rm rad/s"

"v=22.0\: \rm m/s"

"t=9.00\: \rm s"

"d=58.0\:\rm cm"

The final angular speed


"\omega=\frac{v}{R}=\frac{22.0}{0.58/2}=75.9\:\rm rad/s"


(a)


"\phi=\frac{\omega+\omega_0}{2}=\rm \frac{75.9+0.00}{2}\times 9.00=341\: rad""N=\frac{\phi}{2\pi}=\frac{341}{2\pi}=54.4"

(b)


"n=2\pi\omega=6.28\times 75.9\:\rm rad/s=477 \:\rm rev/s"